HUBBLE SPACE TELESCOPE DAILY REPORT #1439 PERIOD COVERED: 0000Z (UTC) 07/24/95 - 0000Z (UTC) 07/25/95 Daily Status Report as of 206/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 6119 (Snapshot Survey for Companions of Planetary Nebula Nuclei) The WF/PC-2 was used to make snapshot survey observations of planetary nebulae NGC-6781 and A61 to study their nuclei and to look for possible companions. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Four Sets of WF/PC-2 6188 (WF/PC-2 Cycle 5 Calibration: Dark) The WF/PC-2 was used to make internal dark calibration images. The observations were completed as scheduled, and no problems were reported. 1.3 Completed FGS 6047 (Calibrating the Mass Luminosity Relation at the End of the Main Sequence) The FGS were used to make astrometric observations of star GL494AB as part of a study to determine its mass luminosity relationship. The purpose is to measure the precise parallax and relative orbit of its faint companion. The observations were completed as planned, and no problems were reported. 1.4 Completed FOS/BL 5492 (Detection of the Intergalactic Medium Using Gunn-Peterson HEII304A Absorption Line) The Faint Object Spectrograph (Blue Side) was used to make spectrographic observations of Quasi-Stellar Object 133647 to search for redshifted 304A line (absorption) of He II. The observations were completed as scheduled, and no anomalies were noted. 1.5 Completed WF/PC-2 6254 (GTO Pure Parallel Program - Cycle 5) The WF/PC-2 was used to make some long exposure observations taken in parallel with the prime science observations. These were a random sample of high galactic latitude fields. The observations were executed as scheduled, and no anomalies were reported. 1.6 Completed GHRS/2/1 5159 (Imaging and Spectroscopy of the Nearby QSO2130+099: Cycle 4 Baseline) The High Resolution Spectrograph (Both Sides One and Two) were used to make observations of nearby quasi-stellar object QSO2130+09. The guide star acquisition failed to coarse track on FGS #1/#3 at 205/1420Z during LOS. No stop, search radius, or step limit flags set at AOS. HSTAR 5246 reported that observations 1 through 7 were impacted. 1.7 Completed WF/PC-2 6189 (WF/PC-2 Cycle 5 Calibration: Internal Monitor - Vis Flats) The WF/PC-2 was used to make internal visual flat calibration images. Both photometric and narrow-band filters were used. The observations were executed as planned, and no problems were reported. 1.8 Completed WF/PC-2 5975 (High Spectral Resolution Studies of Z>4 Quasars) The WF/PC-2 was used to make observations of Quasar BR1202-072 which is known to have a redshift of greater the 4.0. The observed surface brightness distribution of the protogalactic material surrounding these QSOs will allow a constraint to be placed on both the QSO and galaxy formation. The observations was executed as scheduled, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 5 Guide Star acquisition fail to coarse track at 205/1420Z. See Operations Note 1.6 Scheduled Reacquisitions : 9 Successful : 6 See operations Note 1.6 2.2 FHST Updates: Scheduled : 17 Successful : 17 2.3 Operations Notes: No operations anomalies were noted during the reporting period. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-